a rectangular pyramid which has a rectangular base with length 6 and width 10, and height of 18
Answer by avani(25)


Volume of a Rectangular Pyramid: V= 1/3 x B x h
l=length=6
w=width=10
B= Area of Base=Formula to Find Base= l x w = 6 x 10 = 60
H=Hight=18
Thus, V= 1/3 x B x h = 1/3 x 60 x 18 = 360 units